OpenStudy (anonymous):

Suppose R '(1) = 14. If f(x) = R(4x), find f ' (1/4)?

OpenStudy (loser66):

what is it?

OpenStudy (anonymous):

?

OpenStudy (loser66):

where is the problem come from? which grade? which course?

OpenStudy (anonymous):

Calculus I

OpenStudy (loser66):

what does R mean?

OpenStudy (anonymous):

nothing just like if it were F

OpenStudy (loser66):

so? f (x) = R (4x) f'(x) = R'(4x) * (4x) '=4 R' (4x) f'(1/4) = 4 R'(4*1/4) = 4 R'(1) , just plug R'(1) in to get f'(1/4)

OpenStudy (anonymous):

Oh okay I forgot to multiply by the derivative of the inside as well thanks!